OpenStudy (anonymous):

can someone please help me with this? its making my head explode because I am not sure if you need to use the method of substitution. what is the point of intersection? y=-2x 2x+3y=-12

OpenStudy (anonymous):

Use it anyway!

OpenStudy (anonymous):

If you don't know how to do it, then use the method you know!

OpenStudy (anonymous):

Math Rule #159: Do what you know works.

OpenStudy (anonymous):

yeah but will that give me my intersection?

OpenStudy (anonymous):

The point of intersection is basically the values of x and y.

OpenStudy (anonymous):

So if \(x=a\) and \(y=b\) this means the point of intersection is \((a,b)\).

OpenStudy (phi):

solving 2 equations like this with algebra is the same thing as finding where two lines intersect.

OpenStudy (anonymous):

oh okay because my answer was 0,-6 and I think that's wrong

OpenStudy (phi):

when x is 0, the first equation gives you y=0 which is not -6 your solution has to work for both equations

OpenStudy (phi):

try again y=-2x 2x+3y=-12 replace y with -2x in the 2nd equation 2x + 3(-2x) = -12 can you simplify this ?

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

x=3

OpenStudy (phi):

are you guessing ? the first thing is simplify 3*-2*x to -6x in 2x + 3(-2x) = -12 to get 2x -6x = -12 now you have 2 x's takeway 6 x's. how many x's do you have ?

OpenStudy (anonymous):

-4x

OpenStudy (phi):

so you have -4x= -12 divide both sides by -4 -4x/-4 = -12/-4 on the left, -4/-4 is 1, and 1x is x x = ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

3

OpenStudy (phi):

minus divided by minus makes a plus -12/-4 is +3 x= 3 now find y using y=-2x (replace the x with 3 to find y )

OpenStudy (anonymous):

your really good thank you very much

OpenStudy (phi):

yw